please solve

2x-7 over 3 = 2x+3 over 5

thank you

Answer by user_dude2008(1862)

(2x-7)/3 = (2x+3)/5


15[(2x-7)/3] = 15[(2x+3)/5]


5(2x-7) = 3(2x+3)


10x-35 = 6x+9

10x-6x = 9 + 35

4x = 44

x=44/4

x=11


Ans: x = 11